# MergeNet: Knowledge Migration across Heterogeneous Models, Tasks, and Modalities

### v2

@@ Abstract
MergeNet: Knowledge Migration across Heterogeneous Models, Tasks, and Modalities
Kunxi Lithanks: Equal Contributions., Tianyu Zhan11footnotemark: 1, Kairui Fu11footnotemark: 1, Shengyu Zhangthanks: Corresponding author., Kun Kuang, Jiwei Li, Zhou Zhao  and Fei Wu Zhejiang University {kunxili, 3200105641, fukairui.fkr, sy_zhang, kunkuang, jiwei_li, zhaozhou, wufei}@zju.edu.cn
1 Introduction
@@
In an era where edge computing devices are ubiquitous, the deployment of deep neural networks (DNNs) on such devices is often constrained by limited computational resources and storage capacity. This limitation typically necessitates the utilization of smaller neural network architectures, which, while computationally economical, often compromise performance. A promising approach to mitigate this limitation involves the strategic transfer of knowledge from larger, more capable models to these constrained counterparts. A prime example of this approach is knowledge distillation (KD) Hinton et al. (2015); Beyer et al. (2022); Shen et al. (2021). This technique involves training a compact student model to mimic the output logic or intermediate layer representations of a more comprehensive teacher model, thereby improving its accuracy. Transfer learning Zhuang et al. (2020); He et al. (2021); HassanPour Zonoozi and Seydi (2023) offers another approach, commonly through the pre-training and fine-tuning workflow, where knowledge learned during pre-training on large-scale datasets is applied to downstream tasks via backbone sharing. However, these methods typically require similar task types or model architectures, which can limit their applicability. For instance, in diverse Internet of Things (IoT) applications, devices often face varying computational resources and tasks, making it difficult to find matching models for knowledge transfer. Therefore, facilitating heterogeneous knowledge transfer across different model architectures, tasks, and modalities remains a pressing challenge that needs to be addressed.
Unlike previous methods, we consider knowledge transfer between models from a different perspective. We focus on the inherent properties of model parameters, regarding them as the natural carriers of knowledge. An intuitive method is to adopt the idea of parameter sharing, similar to Cai et al. (2021) by putting the smaller model into the larger one, making it a subset of the larger model for additional supervision. We test the effectiveness of parameter sharing on two models with different architectures, ResNet50 He et al. (2016) and MobileNetV2 Sandler et al. (2018), on the CIFAR-100 Krizhevsky et al. (2009). As shown in Figure 1(d), the performance of MobileNetV2 suggests that direct parameter sharing may not be the optimal conduit for heterogeneous knowledge transfer. We speculate that the reasons might be: (i) Direct parameter sharing might disrupt the knowledge of the original modules when heterogeneous modules have significantly different functionalities, like sharing between linear layers and attention mechanism modules. (ii) Typically, larger models contain more advanced knowledge (e.g., fine-grained feature combination) than smaller ones, which the latter might not directly comprehend, leading to potential incompatibility in knowledge between models due to direct parameter sharing.
@@
2 Related Work
In the evolving field of artificial intelligence, knowledge transfer stands out as a crucial mechanism, facilitating the application of insights from specific tasks, domains, or models to new and related contexts. Transfer learning Zhuang et al. (2020); He et al. (2021) is one of the most prevalent forms, involving the extraction of knowledge from one task and applying it to a different but related task to boost the model’s generalization and learning efficiency. This often entails pre-training a model on extensive datasets or complex tasks, followed by fine-tuning it for more specific tasks. Domain adaptation Farahani et al. (2021); HassanPour Zonoozi and Seydi (2023); Ding et al. (2023); Kim et al. (2021) focuses on adapting models trained on one or multiple source domains to function effectively in a distinct target domain. Multi-task learning Crawshaw (2020); Standley et al. (2020); Kurin et al. (2022); Groenendijk et al. (2021) trains models to solve multiple related tasks simultaneously, promoting cross-task information flow through shared representations. This approach allows each task to benefit from features and patterns learned from other tasks. These cross-task learning methods not only improve data utilization and learning efficiency but also enhance the adaptability and robustness of models when confronted with new tasks.
Knowledge distillation (KD) Hinton et al. (2015); Beyer et al. (2022); Shen et al. (2021) is a key knowledge transfer method that involves transferring knowledge from large teacher networks to smaller student models. This is achieved by training the student model to replicate the output logic Hinton et al. (2015); Zhao et al. (2022); Yang et al. (2021); Zhou et al. (2021) or the intermediate layer features Zagoruyko and Komodakis (2016); Yang et al. (2022a) of the teacher model. Recent studies have focused on optimizing the distillation process by improving loss functions and strategies to boost the student’s performance. For instance, MGD Yang et al. (2022b) enhances the transfer process by masking student model features, forcing it to reconstruct the teacher’s characteristic features for a more accurate and robust knowledge transfer. Similarly, NKD Yang et al. (2023) enhances the student model’s use of soft labels by normalizing non-target logits, allowing the student to better utilize the rich information embedded in these labels.
@@ 2 Related Work
3 Method
3.1 Problem Formulation
@@ 3.2 Implementation
Similar to LoRA Hu et al. (2021), we re-encode the parameter matrices $W_{l,t}$ and $W_{s,t}$ involved in knowledge transfer through low-rank decomposition:
@@
$\begin{split}\tilde{A}_{s,t}&=\sum_{i=1}^{4}\omega_{i}\text{Attn}(\phi_{R/L}(A_{s,t}),\\ &\quad\quad\quad\quad\quad\phi_{R/L}(A_{l,t}),\phi_{R/L}(A_{l,t})).\end{split}$ (6)
@@ 4 Experiments
4.1 Cross-Structural Knowledge Transfer
@@
We conduct cross-structure knowledge transfer using the CIFAR-100 Krizhevsky et al. (2009) dataset. This dataset comprises 100 categories, with training and validation sets containing 50k and 10k images, respectively. We use ResNet50 He et al. (2016) as the larger model and MobileNetV2 Sandler et al. (2018) as the smaller model, both of which are not pre-trained. The goal of cross-structure knowledge transfer is to transfer knowledge from one model’s module to a structurally different module in another model. Specifically, we focus on transferring knowledge from the linear classifier of one model to the convolutional layer of another model.
@@ Methods
KD Hinton et al. (2015) 64.32 88.62
RKD Park et al. (2019) 65.48 88.9
DKD Zhao et al. (2022) 65.23 89.01
NKD Yang et al. (2023) 65.09 88.9
@@
KDHinton et al. (2015) 68.36 89.9
RKDPark et al. (2019) 68.6 90.21
DKD Zhao et al. (2022) 69.03 90.25
NKDYang et al. (2023) 69.27 90.18
@@
4.2 Cross-Modal Knowledge Transfer
@@
We conduct experiments in cross-modal knowledge transfer on two distinct tasks: Visual Question Answering using the VQA v2.0 Goyal et al. (2017) dataset and Image-Text Retrieval using the MSCOCO Lin et al. (2014) dataset, with X-VLM Zeng et al. (2021) as the experimental model. For the Image-Text Retrieval task, we use R@10 as the evaluation metric. Given the large size of the datasets and limited computational resources, which led to lengthy training times, we train the model using only 10% of the training set and assess the effectiveness of cross-modal knowledge transfer on the complete test set. Our objective is to use knowledge from one modality to guide the learning of another modality. Specifically, we explore transferring knowledge from the parameters of the visual encoder to the textual encoder and vice versa.
@@
4.3 Cross-Task Knowledge Transfer
@@
The results of cross-task knowledge transfer on the SQuAD v2.0 and IMDb datasets are shown in Table 3. Our method achieve performance improvements in both tasks. For example, transferring knowledge from the classification task to the question-answering task results in a 1.72% increase in Exact Match (EM) score and a 2.37% increase in its F1 score. Similarly, transferring knowledge from the question-answering task to the classification task leads to a 0.52% reduction in error rate. We believe that in similar tasks, the forms of knowledge representation are likely similar, and models performing different tasks can enhance their own task performance by learning the knowledge from other tasks. To demonstrate the capability of our method in aligning knowledge across different tasks, we present a more challenging cross-task knowledge transfer scenario in Appendix A.
@@
Tf-KD Yuan et al. (2020) 65.43 88.56
USKD Yang et al. (2023) 65.66 86.61
@@
4.4 Self Knowledge Transfer.
@@
To thoroughly evaluate the broad applicability of our method, we conduct a series of self knowledge transfer experiments similar to self-distillation on the CIFAR-100 dataset using MobilenetV2. Specifically, we attempt to transfer knowledge from the linear classifier to the 4th, 8th, 12th, and 16th Inverted Residual Blocks (IRB) out of a total of 17, to test the self-knowledge transfer capability of our method. Furthermore, we compare our approach with state-of-the-art self-distillation methods, including Tf-KD Yuan et al. (2020) and USKD Yang et al. (2023). These methods obtain additional supervisory signals by setting manual soft labels.
@@
4.5 Knowledge Transfer from Frozen, Pre-Trained Model
We attempt to transfer the knowledge from a frozen, pre-trained larger model to a smaller one. In the setup, the larger model only needs to retain the parameters involved in the knowledge transfer and does not require forward propagation, resulting in fewer computational requirements. Following the settings in Section 4.1, ResNet50 serves as the pre-trained model. We provide detailed information on the total parameter counts required for each method, denoted as ’Params’. As shown in Table 5, our approach achieves better performance with lower memory overhead.
4.6 Knowledge Transfer Across Different Layers
We explore the effectiveness of our method in transferring knowledge across different layers of information. Similar to Section 4.3, we use BERT and DistilBERT for our experiments. The results, as shown in Table 6, indicate that the most significant performance improvement occurs when both models select their last layers for knowledge transfer. It is widely recognized that deeper layers in neural networks contain more advanced semantic information, enabling a better understanding of higher-level concepts. Therefore, transferring knowledge from these deeper layers imparts richer semantic information to DistilBERT. However, an exception is noted: the transfer between the 6th layer of BERT and the 3rd layer of DistilBERT outperforms the transfer between the 12th layer of BERT and the 3rd layer of DistilBERT. This suggests that while higher-level information typically enhances performance, overly advanced knowledge may be challenging for the model to comprehend, potentially disrupting its existing knowledge structure.
4.7 Ablation Study
@@
We study the impact of the knowledge transfer cycle $T_{cycle}$, which controls the proportion of self-learning during the training process. Figure 3 shows the experimental results of knowledge transfer for MobileNetV2 and ResNet50 on the CIFAR-100 dataset under varying $T_{cycle}$ coefficients. We observe that incorporating self-learning during the training process leads to performance improvements. For example, MobileNetV2 achieves a performance of 66.51% when $T_{cycle}$ is set to 2, representing a 6.42% increase from the 60.09% performance without self-learning. Additionally, higher self-learning ratios lead to a decline in performance for both models. This is because less frequent knowledge transfer hampers the training of the parameter adapter, reducing the effectiveness of the knowledge transfer. Notably, ResNet50 is less affected by changes in the self-learning ratio compared to MobileNetV2, indicating that larger models are better at absorbing knowledge from smaller models.
@@
We conduct an ablation study, as shown in Table 7, to demonstrate the effectiveness of each component in MergeNet. As mentioned in Section 3.2, the parameter adapter can alternatively use MLP as the backbone. In this case, the target model directly adopts the knowledge from the source model while ignoring its own accumulated knowledge, potentially leading to training instability. We compare MergeNet with the MLP-based parameter adapter (row 1 vs row 4). The results show that the MLP-based parameter adapter significantly decreases performance by 1.82% and 1.31% for both models, respectively. Furthermore, we compare MergeNet with different variants of LPKA: (i) LPKA(Individual Attn), which uses only the row-flattened form of the low-rank matrix; (ii) LPKA(Avg Attn), which averages each softmax attention module without using trainable weight parameters. The results indicate that using only the row-flattened form is less effective than considering both row and column, suggesting that column vectors in the parameter matrix also contain crucial model knowledge. Additionally, the performance of MergeNet may decline without trainable weight parameters. These results validate the contribution of each component to enhancing model performance.
5 Conclusion
@@
Limitations
We emphasize that the knowledge embedded in large models can substantially enhance the performance of smaller models. However, identifying meaningful knowledge from the smaller model to benefit the larger one remains a current challenge, worthy of further research efforts.
Appendix A More Challenging Cross-Task Knowledge
Implementation Details.
[…]
Integrated Knowledge Transfer Results.
We conduct cross-structure modal task knowledge transfer experiments, which can be seen as a more challenging form of cross-task knowledge transfer. We choose two significantly different tasks for our experiments: question-answering and image classification. As shown in Table 8, our method is effective in transferring and applying knowledge learned in one task to another, significantly different task. For instance, for MobileNetV2, there is a 2.09% improvement in Top-1 accuracy, and for BERT, there is a 1.79% increase in the F1 score. We believe that despite the significant differences between the tasks used in our experiments, there may be some common information processing mechanisms shared among different tasks. By learning knowledge relevant to their own tasks from other tasks, models can improve their performance.
Appendix B Hyperparameters Used in Experiments
Cross-Structural Knowledge Transfer.
All models are trained using the AdamW optimizer, with a uniform batch size of 128. For the architectures ResNet50 and MobileNetV2, we set the learning rate at 1e-2 and the weight decay at 1e-4. In contrast, for MergeNet, a more intricate setup is adopted: the learning rate is reduced to 1e-3, and the weight decay is increased to 1e-2. Furthermore, MergeNet is configured with a depth of 2 layers, and the LPKA is equipped with 8 softmax attention heads in each layer. Notably, the knowledge transfer cycle $T_{cycle}$ is meticulously calibrated for optimal efficacy, set at 4 for transfers from ResNet50 to MobileNetV2, and 5 for the reverse direction, ensuring a robust and systematic knowledge distillation process.
Cross-Modal Knowledge Transfer.
For the VQA task, we tailor the settings to accommodate the specific demands of the task: a batch size of 16, a base model learning rate of 5e-4, and a weight decay of 1e-2. Similarly, for the ITR task, the batch size is maintained at 16, while the base model learning rate is set slightly lower at 1e-4, accompanied by a weight decay of 1e-2, aligning with the task’s distinct characteristics. Moreover, MergeNet is carefully structured with 2 layers, and the LPKA is configured with 12 softmax attention heads per layer, ensuring a rich and attentive knowledge integration. The knowledge transfer cycle $T_{cycle}$ is concisely set at 2.
Methods CIFAR-100 SQuAD v2.0
Top-1 Acc Top-5 Acc EM F1
Vanilla 63.87 88.77 70.17 73.06
MergeNet 65.56 88.74 70.89 74.15
Cross-Task Knowledge Transfer.
All models are trained using the AdamW optimizer. For the question-answering task, we meticulously calibrat the training parameters, setting the batch size at 32, the base model learning rate at 5e-5, and the weight decay at a fine-tuned 1e-5. Similarly, for the classification task, we adopt a batch size of 6, maintaining the base model learning rate at 5e-5, while adjusting the weight decay to 1e-2 to cater to the task’s specific demands. Furthermore, MergeNet is architected with 2 layers, and the LPKA is enriched with 12 softmax attention heads per layer. The knowledge transfer cycle $T_{cycle}$ is meticulously calibrated for optimal efficacy, set at 1 for transfers from the question-answering task to the classification task, and 2 for the reverse direction, ensuring a robust and systematic knowledge distillation process. Notably, in More Challenging Cross-Task Knowledge experiments, the hyperparameter settings for the models align consistently with the corresponding configurations delineated above.
Self Knowledge Transfer.
We train all models using the AdamW optimizer. For the base models, their configurations are the same as those used in cross-structure knowledge transfer. Turning our focus to MergeNet, we calibrate its learning rate to 1e-3 and set the weight decay to 1e-2, with the knowledge transfer cycles $T_{cycle}$ uniformly set to 4. Architecturally, MergeNet is designed with a depth of 2 layers. Furthermore, LPKA within MergeNet is configured with 8 adept softmax attention heads per layer.

### v3

@@ Abstract
Introduction
@@
In an era where edge computing devices are ubiquitous, the deployment of deep neural networks (DNNs) on such devices is often constrained by limited computational resources and storage capacity. This limitation typically necessitates the utilization of smaller neural network architectures, which, while computationally economical, often compromise performance. A promising approach to mitigate this limitation involves the strategic transfer of knowledge from larger, more capable models to these constrained counterparts. A prime example of this approach is knowledge distillation (KD) (Hinton et al. 2015; Beyer et al. 2022; Shen et al. 2021). This technique involves training a compact student model to mimic the output logic or intermediate layer representations of a more comprehensive teacher model, thereby improving its accuracy. Transfer learning (Zhuang et al. 2020; He et al. 2021; HassanPour Zonoozi and Seydi 2023) offers another approach, commonly through the pre-training and fine-tuning workflow, where knowledge learned during pre-training on large-scale datasets is applied to downstream tasks via backbone sharing. However, these methods typically require similar task types or model architectures, which can limit their applicability. For instance, in diverse Internet of Things (IoT) applications, devices often face varying computational resources and tasks, making it difficult to find matching models for knowledge transfer. Therefore, facilitating heterogeneous knowledge transfer across different model architectures, tasks, and modalities remains a pressing challenge that needs to be addressed.
Unlike previous methods, we consider knowledge transfer between models from a different perspective. We focus on the inherent properties of model parameters, regarding them as the natural carriers of knowledge. An intuitive method is to adopt the idea of parameter sharing, similar to Cai et al. (2021) by putting the smaller model into the larger one, making it a subset of the larger model for additional supervision. We test the effectiveness of parameter sharing on two models with different architectures, ResNet50 (He et al. 2016) and MobileNetV2 (Sandler et al. 2018), on the CIFAR-100 (Krizhevsky, Hinton et al. 2009). As shown in Figure 1(d), the performance of MobileNetV2 suggests that direct parameter sharing may not be the optimal conduit for heterogeneous knowledge transfer. We speculate that the reasons might be: (i) Direct parameter sharing might disrupt the knowledge of the original modules when heterogeneous modules have significantly different functionalities, like sharing between linear layers and attention mechanism modules. (ii) Typically, larger models contain more advanced knowledge (e.g., fine-grained feature combination) than smaller ones, which the latter might not directly comprehend, leading to potential incompatibility in knowledge between models due to direct parameter sharing.
@@
Related Work
In the evolving field of artificial intelligence, knowledge transfer stands out as a crucial mechanism, facilitating the application of insights from specific tasks, domains, or models to new and related contexts (Fu et al. 2024). Transfer learning (Zhuang et al. 2020; He et al. 2021) is one of the most prevalent forms, involving the extraction of knowledge from one task and applying it to a different but related task to boost the model’s generalization and learning efficiency. This often entails pre-training a model on extensive datasets or complex tasks, followed by fine-tuning it for more specific tasks. Domain adaptation (Farahani et al. 2021; HassanPour Zonoozi and Seydi 2023; Ding et al. 2023; Kim et al. 2021) focuses on adapting models trained on one or multiple source domains to function effectively in a distinct target domain. Multi-task learning (Crawshaw 2020; Standley et al. 2020; Kurin et al. 2022; Groenendijk et al. 2021) trains models to solve multiple related tasks simultaneously, promoting cross-task information flow through shared representations. This approach allows each task to benefit from features and patterns learned from other tasks. These cross-task learning methods not only improve data utilization and learning efficiency but also enhance the adaptability and robustness of models when confronted with new tasks.
Knowledge distillation (KD) (Hinton et al. 2015; Beyer et al. 2022; Shen et al. 2021) is a key knowledge transfer method that involves transferring knowledge from large teacher networks to smaller student models. This is achieved by training the student model to replicate the output logic (Hinton et al. 2015; Zhao et al. 2022; Yang et al. 2021; Zhou et al. 2021) or the intermediate layer features (Zagoruyko and Komodakis 2016; Yang et al. 2022a) of the teacher model. Recent studies have focused on optimizing the distillation process by improving loss functions and strategies to boost the student’s performance. For instance, MGD (Yang et al. 2022b) enhances the transfer process by masking student model features, forcing it to reconstruct the teacher’s characteristic features for a more accurate and robust knowledge transfer. Similarly, NKD (Yang et al. 2023) enhances the student model’s use of soft labels by normalizing non-target logits, allowing the student to better utilize the rich information embedded in these labels.
@@
Method
Problem Formulation
@@
Similar to LoRA (Hu et al. 2021), we re-encode the parameter matrices $W_{l,t}$ and $W_{s,t}$ involved in knowledge transfer through low-rank decomposition:
@@
$\begin{split}\tilde{A}_{s,t}&=\sum_{i=1}^{4}\omega_{i}\text{Attn}(\phi_{R/L}(A_{s,t}),\phi_{R/L}(A_{l,t}),\phi_{R/L}(A_{l,t})).\end{split}$ (6)
@@
Cross-Structural Knowledge Transfer
@@
We conduct cross-structure knowledge transfer using the CIFAR-100 (Krizhevsky, Hinton et al. 2009) dataset. This dataset comprises 100 categories, with training and validation sets containing 50k and 10k images, respectively. We use ResNet50 (He et al. 2016) as the larger model and MobileNetV2 (Sandler et al. 2018) as the smaller model, both of which are not pre-trained. The goal of cross-structure knowledge transfer is to transfer knowledge from one model’s module to a structurally different module in another model. Specifically, we focus on transferring knowledge from the linear classifier of one model to the convolutional layer of another model.
@@ Methods
KD (Hinton et al. 2015) 64.32 88.62
RKD (Park et al. 2019) 65.48 88.9
DKD (Zhao et al. 2022) 65.23 89.01
NKD (Yang et al. 2023) 65.09 88.9
@@
KD(Hinton et al. 2015) 68.36 89.9
RKD(Park et al. 2019) 68.6 90.21
DKD (Zhao et al. 2022) 69.03 90.25
NKD(Yang et al. 2023) 69.27 90.18
@@
Cross-Modal Knowledge Transfer
@@
We conduct experiments in cross-modal knowledge transfer on two distinct tasks: Visual Question Answering using the VQA v2.0 (Goyal et al. 2017) dataset and Image-Text Retrieval using the MSCOCO (Lin et al. 2014) dataset, with X-VLM (Zeng, Zhang, and Li 2021) as the experimental model. For the Image-Text Retrieval task, we use R@10 as the evaluation metric. Given the large size of the datasets and limited computational resources, which led to lengthy training times, we train the model using only 10% of the training set and assess the effectiveness of cross-modal knowledge transfer on the complete test set. Our objective is to use knowledge from one modality to guide the learning of another modality. Specifically, we explore transferring knowledge from the parameters of the visual encoder to the textual encoder and vice versa.
@@
Cross-Task Knowledge Transfer
@@
The results of cross-task knowledge transfer on the SQuAD v2.0 and IMDb datasets are shown in Table 3. Our method achieve performance improvements in both tasks. For example, transferring knowledge from the classification task to the question-answering task results in a 1.72% increase in Exact Match (EM) score and a 2.37% increase in its F1 score. Similarly, transferring knowledge from the question-answering task to the classification task leads to a 0.52% reduction in error rate. We believe that in similar tasks, the forms of knowledge representation are likely similar, and models performing different tasks can enhance their own task performance by learning the knowledge from other tasks. To demonstrate the capability of our method in aligning knowledge across different tasks, we present a more challenging cross-task knowledge transfer scenario in the following section.
Methods CIFAR-100 SQuAD v2.0
Top-1 Acc Top-5 Acc EM F1
Vanilla 63.87 88.77 70.17 73.06
MergeNet 65.56 88.74 70.89 74.15
More Challenging Cross-Task Knowledge
Implementation Details.
[…]
Integrated Knowledge Transfer Results.
We conduct cross-structure modal task knowledge transfer experiments, which can be seen as a more challenging form of cross-task knowledge transfer. We choose two significantly different tasks for our experiments: question-answering and image classification. As shown in Table 4, our method is effective in transferring and applying knowledge learned in one task to another, significantly different task. For instance, for MobileNetV2, there is a 1.69% improvement in Top-1 accuracy, and for BERT, there is a 1.09% increase in the F1 score. We believe that despite the significant differences between the tasks used in our experiments, there may be some common information processing mechanisms shared among different tasks. By learning knowledge relevant to their own tasks from other tasks, models can improve their performance.
@@
Tf-KD (Yuan et al. 2020) 65.43 88.56
USKD (Yang et al. 2023) 65.66 86.61
@@
Self Knowledge Transfer
@@
To thoroughly evaluate the broad applicability of our method, we conduct a series of self knowledge transfer experiments similar to self-distillation on the CIFAR-100 dataset using MobilenetV2. Specifically, we attempt to transfer knowledge from the linear classifier to the 4th, 8th, 12th, and 16th Inverted Residual Blocks (IRB) out of a total of 17, to test the self-knowledge transfer capability of our method. Furthermore, we compare our approach with state-of-the-art self-distillation methods, including Tf-KD (Yuan et al. 2020) and USKD (Yang et al. 2023). These methods obtain additional supervisory signals by setting manual soft labels.
@@
Knowledge Transfer from Frozen, Pre-Trained Model
We attempt to transfer the knowledge from a frozen, pre-trained larger model to a smaller one. In the setup, the larger model only needs to retain the parameters involved in the knowledge transfer and does not require forward propagation, resulting in fewer computational requirements. Following the settings in the previous section, ResNet50 serves as the pre-trained model. We provide detailed information on the total parameter counts required for each method, denoted as ’Params’. As shown in Table 6, our approach achieves better performance with lower memory overhead.
Knowledge Transfer Across Different Layers
We explore the effectiveness of our method in transferring knowledge across different layers of information. Similar to the previous section, we use BERT and DistilBERT for our experiments. The results, as shown in Table 7, indicate that the most significant performance improvement occurs when both models select their last layers for knowledge transfer. It is widely recognized that deeper layers in neural networks contain more advanced semantic information, enabling a better understanding of higher-level concepts. Therefore, transferring knowledge from these deeper layers imparts richer semantic information to DistilBERT. However, an exception is noted: the transfer between the 6th layer of BERT and the 3rd layer of DistilBERT outperforms the transfer between the 12th layer of BERT and the 3rd layer of DistilBERT. This suggests that while higher-level information typically enhances performance, overly advanced knowledge may be challenging for the model to comprehend, potentially disrupting its existing knowledge structure.
Ablation Study
@@
We study the impact of the knowledge transfer cycle $T_{cycle}$, which controls the proportion of self-learning during the training process. Figure 3 shows the experimental results of knowledge transfer for MobileNetV2 and ResNet50 on the CIFAR-100 dataset under varying $T_{cycle}$ coefficients. We observe that incorporating self-learning during the training process leads to performance improvements. For example, MobileNetV2 achieves a performance of 66.51% when $T_{cycle}$ is set to 4, representing a 6.42% increase from the 60.09% performance without self-learning. Additionally, higher self-learning ratios lead to a decline in performance for both models. This is because less frequent knowledge transfer hampers the training of the parameter adapter, reducing the effectiveness of the knowledge transfer. Notably, ResNet50 is less affected by changes in the self-learning ratio compared to MobileNetV2, indicating that larger models are better at absorbing knowledge from smaller models.
@@
We conduct an ablation study, as shown in Table 8, to demonstrate the effectiveness of each component in MergeNet. As mentioned in the previous section, the parameter adapter can alternatively use MLP as the backbone. In this case, the target model directly adopts the knowledge from the source model while ignoring its own accumulated knowledge, potentially leading to training instability. We compare MergeNet with the MLP-based parameter adapter (row 1 vs row 4). The results show that the MLP-based parameter adapter significantly decreases performance by 1.82% and 1.31% for both models, respectively. Furthermore, we compare MergeNet with different variants of LPKA: (i) LPKA(Individual Attn), which uses only the row-flattened form of the low-rank matrix; (ii) LPKA(Avg Attn), which averages each softmax attention module without using trainable weight parameters. The results indicate that using only the row-flattened form is less effective than considering both row and column, suggesting that column vectors in the parameter matrix also contain crucial model knowledge. Additionally, the performance of MergeNet may decline without trainable weight parameters. These results validate the contribution of each component to enhancing model performance.
Conclusion
@@
Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62441605, 62402429), the National Science and Technology Major Project (No. 2022ZD0119100) and the Key Research and Development Program of Zhejiang Province (No. 2024C03270). This work was also partially supported by the ZJU Kunpeng&Ascend Center of Excellence and the Ningbo Yongjiang Talent Introduction Programme (No. 2023A-397-G).